[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jun 29, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jun 29 13:52:52 UTC 2016


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
GFE/NCP issues: TAFB could not publish Atlantic NH2 grids to Official from  
either LX2 or LX6 during the swing shift. Also, Pacific NH1 desk had issues  
sending grids to NAWIPS (PMSL) and had to close and reopen GFE several  
times. NCF traced the problem to the EDEX server; NCF then restarted EDEX  
and GFE performed normally again.


HSU:
Duplicate copies of older shift logs continue to be sent via e-mail. TSB is  
investigating.
Other items:
Monica indicated that NHC will NOT be switching to our local GOES-E ingest.  
NCO and TSB are still investigating data flow and quality control issues.


FYI from NASA Sport:
At 8am CDT, there was a change to the Sport data NHC received in the  
$SAT/GOES_R-PG directory. Forecasters can check NMAP and compare with the  
details listed in Monica's email sent out yesterday.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160629/d440132b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list